# Is ChatGPT Transforming Academics’ Writing Style?

### v1

@@ Abstract
Based on one million arXiv papers submitted from May 2018 to January 2024, we assess the textual density of ChatGPT’s writing style in their abstracts by means of a statistical analysis of word frequency changes. Our model is calibrated and validated on a mixture of real abstracts and ChatGPT-modified abstracts (simulated data) after a careful noise analysis. We find that ChatGPT is having an increasing impact on arXiv abstracts, especially in the field of computer science, where the fraction of ChatGPT-revised abstracts is estimated to be approximately 35%, if we take the output of one of the simplest prompts, “revise the following sentences”, as a baseline. We conclude with an analysis of both positive and negative aspects of the penetration of ChatGPT into academics’ writing style.
ChatGPT
@@ 1 Introduction
Since its official release on November 30, 2022, ChatGPT (Chat Generative Pre-trained Transformer) has impacted many aspects of our lives, and academic writing has not been immune. While ChatGPT does increase productivity and may help scientific discovery (Noy & Zhang, 2023; AI4Science & Quantum, 2023), we must be wary of its potential risks and the possibility of negative impacts. A large number of papers have already explored the advantages and disadvantages of large language models (LLMs) (Kasneci et al., 2023); here, we focus on ChatGPT, which is being very widely used (reaching an unprecedented 100 million active users three months after its release) and is considered to be one of the two major milestones of language models along with GPT-4. (Zhao et al., 2023; von Garrel & Mayer, 2023)
@@
While there is already a corpus of current research on using ChatGPT in academia (Casal & Kessler, 2023; Lingard et al., 2023; Fergus et al., 2023; Lund et al., 2023), to our knowledge only a handful of works have attempted to quantify its impact on the whole academic community. As this article was being finalized, two preprints appeared that addressed related questions: one focuses on AI conferences peer reviews (Liang et al., 2024a) and, even more recently, analyzes scientific papers(Liang et al., 2024b). They claim that the usage of LLMs is evident in AI conference reviews and scientific writings, especially in computer science papers. Within the broad field of academic writing and publishing, we chose the abstracts of articles as the focus of this work, as they have a relatively uniform format across disciplines, are supposed to condense an entire research article and thus are often highly polished, and can be considered short articles of pure text, not involving pictures nor tables.
@@
ChatGPT is of course able to generate abstracts directly given a suitable prompt (Luo et al., 2023), and studies have shown that identifying such abstracts is not easy even if they remain unedited by humans (Gao et al., 2023; Cheng et al., 2023) – watermarking being a possible strategy to enable such identification (Kirchenbauer et al., 2023). Determining whether a given few sentences were generated by ChatGPT is difficult, but determining that millions of sentences were influenced by ChatGPT is statistically feasible, as we demonstrate here. We analyzed the fingerprints of ChatGPT on scientific abstracts as a function of time in order to tease out a statistical signature, rather than attempting to detect whether a given abstract was generated or polished with ChatGPT.
@@
In fact, that the abstract of a paper shows what we call the “ChatGPT style” does not necessarily mean that the authors directly utilized ChatGPT to generate or modify it. It is also possible that the authors used ChatGPT in another context and that, as a result, their writing habits were influenced by the ChatGPT style – not a remote possibility. It is worth considering in this context that reading and writing in English is more difficult for non-native English academics (Amano et al., 2023). Before ChatGPT was released, the pros and cons of other tools were discussed, such as Google Translate (Mundt & Groves, 2016) and Grammarly (Fitria, 2021), but ChatGPT has a much wider range of application scenarios – not to mention, a much higher flexibility.
@@
We have seen similar AI-induced seismic shifts in the past: after AlphaGo (Silver et al., 2017) shocked the world, professional Go players have begun training with AI, and the sport of Go has been profoundly changed as a result (Kang et al., 2022). AlphaFold brings new opportunities for life science research (Varadi & Velankar, 2023) and ChatGPT has also been used for data extraction in materials science. (Polak & Morgan, 2023). A similar story may be happening with academic writing, especially for researchers whose first language is not English (Hwang et al., 2023). This paper is a first effort at establishing whether this is the case.
@@ 2 Data
arXiv dataset: The metadata of arXiv papers are provided by Kaggle (arXiv.org submitters, 2024). Because the abstracts in this dataset are updated when authors submit changes, we used the first version in 2024 (version 161) as well as the last version before the ChatGPT era (version 105). Our observations and analysis are based on one million arXiv articles submitted from May 2018 to January 2024.
English word frequency: Google Ngram dataset is chosen for comparison and reference (Michel et al., 2011). Specifically, we used the freely available mirrors on Kaggle (http://kaggle.com/datasets/wheelercode/english-word-frequency-list) covering word frequencies from the 1800s to 2019 as established from Google Books.
@@ 3.1 Changes in word frequency
We approach the problem by analyzing how the frequency of words changes after ChatGPT has been deployed, so we define the change factor in the frequency of word $i$, $R_{i}$, as follows:
$R_{i}=\frac{\max_{t}(f_{i}(t))-\min_{t}(f_{i}(t))}{\max_{t}(f_{i}(t))}$ (1)
where $f_{i}(t)$ is the count of word $i$ during the time period $t$. We divided the 1 million abstracts into 100 uneven time-periods, each encompassing 10,000 abstracts.
Figure 1: The 12 words with the highest change rate $R_{i}$ and satisfying $\max_{t}(f_{i}(t))>500$. The vertical red dashed line demarcates the first time period after ChatGPT’s release.
Figure 1 illustrates that most of the words with the largest change rate in the time period considered (generally, an increase) in the abstracts are related to hot research topics of the last few years, such as “Covid-19”, “LLMs”, “AI”. But this is not the case for all words with the largest growth in terms of frequency. Indeed, the frequency of some non-specialized words also starts to skyrocket in early 2023, as presented in Figure 2. How could the frequencies of words like “significant” grow significantly together?
Figure 2: Examples of words with rapidly growing frequency in arXiv abstracts.
Figure 3: The words “are” and “is” are decreasing in frequency in arXiv abstracts.
Another striking example is the frequency change of the words “are” and “is”, as depicted in Figure 3. The counts in 10,000 abstracts of these two words were quite stable before 2023. However, the frequency of these two terms has dropped by more than 10% in 2023.
Because the average length of abstracts tends to grow over time, we also considered normalizing frequencies to abstract length. The corresponding figures are displayed in the Appendix, and show similar trends.
These examples, anecdotal as they are, may represent the tip of the iceberg of a wider and growing phenomenon: the rapid increase in the usage of ChatGPT. The rise and fall in frequency of specific technical nouns may well be related to the changing popularity of certain research topics, but that a research trend is responsible for the change in usage of adjectives appears implausible – even less so for words like “is” and “are”.
3.2 ChatGPT simulations
We wanted to be more specific about the impact of ChatGPT on articles from different disciplines, so we examined arXiv abstracts from different categories separately. The one million arXiv articles were divided into 20 periods in this part in order to increase the number of articles per period and reduce estimation error, which is not the same as the previous part. The identifier numbers of the first and last arXiv articles corresponding to each period are given in the Appendix.
@@ 3.2 ChatGPT simulations
where $q^{d}_{ij}$ represents the word frequency of real abstracts in the dataset and $\tilde{q}^{d}_{ij}$ means the frequency after ChatGPT processing. We have no way of knowing the real usage scenarios of ChatGPT, so some simple prompts were used, for example,
@@
GPT-3.5 was utilized in our simulations for 10,000 abstracts in period 14 (April 2022 to July 2022), although it may have different word preferences than the more recent GPT-4. Many words have different frequencies before and after ChatGPT processing, such as the words “is”, “are”, and “significant” that we mentioned earlier. For simplicity, the results of the 4 categories with the highest number of articles are shown in Table 1 and the rest parts in this paper, namely cs (computer science), math (mathematics), astro (astrophysics), and cond-mat (condensed matter).
Table 1: Word frequency (per abstract) before and after ChatGPT processing in simulation data of period 14.
word category before after change
is cs 2.01 1.73 -14%
is math 1.78 1.61 -9%
is astro 2.13 1.90 -11%
is cond-mat 2.00 1.68 -16%
are cs 1.00 0.83 -17%
are math 0.74 0.71 -5%
are astro 1.39 1.25 -1%
are cond-mat 0.92 0.80 -13%
@@
This corroborates the hypothesis, formulated earlier, that the drop in the frequency of these two words observed in real abstracts in 2023 may have been caused by ChatGPT.
In the meantime, we also defined the word frequency change in all abstracts from year $t-1$ to year $t$, $R_{ij,t}$:
$R_{ij,t}=\frac{F_{ij,t}-F_{ij,t-1}}{F_{ij,t-1}}\,,$ (3)
Only words with a frequency larger than 0.1 times per abstract before ChatGPT processing are plotted in Figure 4 and Figure 5. The correlation coefficient between the word frequency change in arXiv abstracts and our estimated ChatGPT-induced word frequency change is very small in all four categories of abstracts, as shown in Figure 4.
Figure 4: Comparison of the predicted frequency change rate due to ChatGPT $\hat{r}_{ij}$ (x-axis) and the actual word frequency change for all abstracts from 2021 to 2022 $R_{ij,2022}$ (y-axis). CC indicates the correlation coefficient.
Figure 5: Comparison of the predicted frequency change rate due to ChatGPT $\hat{r}_{ij}$ (x-axis) and the actual word frequency change for all abstracts from 2022 to 2023 $R_{ij,2023}$ (y-axis). The correlation coefficients (CC) are now significantly positive.
However, Figure 5 presents a totally different pattern, where $\hat{r}_{ij}$ and $R_{ij,2023}$ are strongly correlated, especially in computer science abstracts. Although many words seem insensitive to ChatGPT, we can still see a positive correlation for some words in this figure, even among the other categories.
Taken together, our consideration point to ChatGPT as one of the important reasons, possibly even the main reason, for the recent word frequency change in abstracts. Our next step is to start by modeling ChatGPT impact, as well as estimating the impact based on real data and simulations.
4 Models and methods
4.1 ChatGPT impact
Imagine different scenarios of using ChatGPT in scientific writing: a researcher might simply use it to correct grammatical errors, another employs it for translating native sentences into English, and yet another one wants it to polish their draft in English very purposefully. In theory, each of these use cases contributes the same proportion of ChatGPT usage. But, as is well known, different prompts will lead to different outputs, which means different word frequency changes. Therefore, we use the more neutral term “ChatGPT impact” instead of “proportion” in our estimation part.
@@ 4.1 ChatGPT impact
$\begin{split}\bar{f}_{ij}(t)=&(1-\eta_{j}(t))f^{*}_{ij}(t)+\eta_{j}(t)f^{*}_{ij}(t)(\bar{r}_{ij}+1)\\ =&f^{*}_{ij}(t)+\eta_{j}(t)f^{*}_{ij}(t)\bar{r}_{ij}\end{split}$ (5)
where $\eta_{j}(t)$ denotes the proportion of abstracts in category $j$ affected by ChatGPT, and $f^{*}_{ij}(t)$ represents the original evolution in word frequency without ChatGPT.
This model is highly idealized: we have to additionally consider the effects of noise (such as randomness inside ChatGPT), uncertainty in word usage evolution without ChatGPT, and the epistemic uncertainty in how users actually prompt ChatGPT.
@@ 4.2 Noise model
We now consider the noise terms, which might be modelled in many different ways.
For example, we denote the word frequency for word $i$ in category $j$ by ${f}^{d}_{ij}$, which represents the word frequency observed in the data:
@@
where $\delta_{ij}(\cdot)$ represents noise and word usage variability which are not directly related to the internal parameters of ChatGPT.
Thus, we can define
$f^{\delta,\eta}_{ij}(t)=\eta_{j}(t)f^{*}_{ij}(t)+\delta_{ij}(\eta_{j}(t)f^{*}_{ij}(t))\,,$ (7)
and
$\begin{split}f^{\delta,1-\eta}_{ij}(t)=&(1-\eta_{j}(t))f^{*}_{ij}(t)\\ &+\delta_{ij}((1-\eta_{j}(t))f^{*}_{ij}(t))\,.\end{split}$ (8)
@@
$f^{d}_{ij}(t)=f^{\delta,1-\eta}_{ij}(t)+\mathrm{C}_{ij}(f^{\delta,\eta}_{ij}(t))$ (9)
where the function $\mathrm{C}_{ij}(\cdot)$ means the frequency after ChatGPT process.
We assume that the noise for word $i$ due to ChatGPT processing can be represented as $\epsilon_{ij}(\cdot)$ and $\epsilon_{ij}^{s}(\cdot)$, then Eq. (2) and Eq. (4) are related by
@@
Then, Eq. (9) – representing the difference in word frequency before and after ChatGPT processing – can be rewritten as
@@
$\begin{split}\xi_{ij}(t)=&(\hat{r}_{ij}+1+\epsilon^{\eta}_{ij}(q,f,t))\delta_{ij}(\eta_{j}(t)f^{*}_{ij}(t))\\ &+\delta^{\prime}_{ij}((1-\eta_{j}(t))f^{*}_{ij}(t))\,.\end{split}$ (16)
where $\delta^{\prime}_{ij}(\cdot)$ follows the same distribution as $\delta_{ij}(\cdot)$. It should be noted that $g_{ij}(t)$ includes only ChatGPT-related noise $\epsilon_{ij}(\cdot)$ and $\epsilon^{s}_{ij}(\cdot)$, however $\xi_{ij}(t)$ contains $\delta_{ij}(\cdot)$ and $\delta^{\prime}_{ij}(\cdot)$ that are unrelated to ChatGPT.

### v2

@@ Abstract
Based on one million arXiv papers submitted from May 2018 to January 2024, we assess the textual density of ChatGPT’s writing style in their abstracts through a statistical analysis of word frequency changes. Our model is calibrated and validated on a mixture of real abstracts and ChatGPT-modified abstracts (simulated data) after a careful noise analysis. The words used for estimation are not fixed but adaptive, including those with decreasing frequency. We find that large language models (LLMs), represented by ChatGPT, are having an increasing impact on arXiv abstracts, especially in the field of computer science, where the fraction of LLM-style abstracts is estimated to be approximately 35%, if we take the responses of GPT-3.5 to one simple prompt, “revise the following sentences”, as a baseline. We conclude with an analysis of both positive and negative aspects of the penetration of LLMs into academics’ writing style.
@@ 1 Introduction
Since ChatGPT (Chat Generative Pre-trained Transformer) was released on November 30, 2022, large language models (LLMs) have started to be widely exploited and therefore affect many aspects of our lives. In this paper, we are concerned with whether LLMs, represented by ChatGPT, are transforming academic writing.
Many papers have already explored the advantages and disadvantages of LLMs (Kasneci et al., 2023). Although they can increase productivity and may help scientific discovery (Noy & Zhang, 2023; AI4Science & Quantum, 2023), the potential risks of using LLMs in academia cannot be ignored (Lund et al., 2023), such as generating incorrect references (Walters & Wilder, 2023).
Researchers have been working on machine-generated text detection since a few years ago (Bakhtin et al., 2019; Gehrmann et al., 2019), which have attracted more attention shortly after ChatGPT appeared (Mitchell et al., 2023; Guo et al., 2023). At the same time, questions have been raised about the reliability of these detectors (Sadasivan et al., 2023). Detection and counter-detection of LLM-generated text soon developed cat-and-mouse games, such as watermarking (Kirchenbauer et al., 2023), paraphrasing (Sadasivan et al., 2023), and the combination of both (Krishna et al., 2024). Besides, distinguishing between human and LLM writing samples is also sometimes not easy for human experts (Casal & Kessler, 2023).
@@
While there is already a corpus of current research on using ChatGPT in academia (Casal & Kessler, 2023; Lingard et al., 2023; Fergus et al., 2023; Lund et al., 2023), to our knowledge only a handful of works have attempted to quantify its impact on the whole academic community. It was only when the first version of this paper was completed that two preprints appeared that addressed related questions: one focuses on AI conferences peer reviews (Liang et al., 2024a) and analyzes scientific papers (Liang et al., 2024b). They claim that the usage of LLMs is evident in AI conference reviews and scientific writings, especially in computer science papers.
Within the broad field of academic writing and publishing, we chose the abstracts of articles as the focus of this work, as they have a relatively uniform format across disciplines, are supposed to condense an entire research article and thus are often highly polished, and can be considered short articles of pure text, not involving pictures nor tables.
@@
Of course, LLMs can generate abstracts directly given a suitable prompt (Luo et al., 2023), and studies have shown that identifying such abstracts is not easy even if they remain unedited by humans  (Gao et al., 2023; Cheng et al., 2023). We have previously discussed methods for detecting LLM-generated text, but the detection of a mixture of human and machine-generated text is usually much harder (Krishna et al., 2024; Zhang et al., 2024). Determining whether a given few sentences were generated by LLMs is difficult, but it is feasible to estimate the extent to which millions of sentences are influenced by LLMs. We analyzed the fingerprints of LLMs on scientific abstracts as a function of time in order to tease out a statistical signature, rather than a binary classification.
@@
In fact, that the abstract of a paper shows what we call the “ChatGPT style” or “LLM style” does not necessarily mean that the authors directly utilized LLM to generate or modify it. It is also possible that the authors used LLM in another context and that, as a result, their writing habits were influenced by the LLM style – not a remote possibility.
It is worth considering in this context that reading and writing in English is more difficult for non-native English academics (Amano et al., 2023). Before ChatGPT was released, the pros and cons of other tools were discussed, such as Google Translate (Mundt & Groves, 2016) and Grammarly (Fitria, 2021), but ChatGPT has a much wider range of application scenarios – not to mention, a much higher flexibility.
@@
We have seen similar AI-induced seismic shifts in the past: after AlphaGo (Silver et al., 2017) shocked the world, professional Go players have begun training with AI, and the sport of Go has been profoundly changed as a result (Kang et al., 2022). A similar story may be happening with academic writing, especially for researchers whose first language is not English (Hwang et al., 2023). This paper is a first effort at establishing whether this is the case.
We also think that analytical rigor is a higher priority than comprehensiveness, and the former is our focus in this paper. Once the reliability of a single analysis is assured, the comprehensive analysis can be more convincing. For example, we should use a more adaptive approach to selecting words for estimation, as well as considering words with decreasing frequency.
@@ 2 Data
arXiv dataset
The metadata of arXiv papers are provided by Kaggle (arXiv.org submitters, 2024). Because the abstracts in this dataset are updated when authors submit changes, we used the first version in 2024 (version 161) as well as the last version before the ChatGPT era (version 105). Our observations and analysis are based on one million arXiv articles submitted from May 2018 to January 2024.
English word frequency
Google Ngram dataset is chosen for comparison and reference  (Michel et al., 2011). Specifically, we used the freely available mirrors on Kaggle (http://kaggle.com/datasets/wheelercode/english-word-frequency-list) covering word frequencies from the 1800s to 2019 as established from Google Books.
@@ 3.1 Changes in word frequency
We approach the problem by analyzing how the frequency of words changes after ChatGPT has been deployed. The frequency of some non-specialized words also starts to skyrocket in early 2023, as presented in Figure 1 (1 million abstracts are divided into 100 uneven time-periods, each encompassing 10,000 abstracts). The number of arXiv articles is getting more and more each month. In general, the larger the sample (the greater the number of articles), the more accurate the estimate will be. We used a similar number of articles in each period rather than the same time interval to keep the error in the estimates on the same scale, providing the same quality of observation and estimation.
(a) Examples of words with rapidly growing frequency in arXiv abstracts.
(b) The words “are” and “is” are decreasing in frequency in arXiv abstracts.
Figure 1: Word frequency changes in abstracts. The vertical red dashed line demarcates the first time period after ChatGPT’s release.
How could the frequencies of words like “significant” grow significantly together? Another striking example is the frequency change of the words “are" and “is”. The counts in 10,000 abstracts of these two words were quite stable before 2023. However, the frequency of these two terms has dropped by more than 10% in 2023.
These examples, anecdotal as they are, may represent the tip of the iceberg of a wider and growing phenomenon: the rapid increase in the usage of ChatGPT or other LLMs. The rise and fall in frequency of specific technical nouns may well be related to the changing popularity of certain research topics, but that a research trend is responsible for the change in usage of adjectives appears implausible – even less so for words like “is” and “are”.
3.2 LLM simulations
We wanted to be more specific about the impact of LLMs on articles from different disciplines, so the arXiv abstracts from different categories were examined separately. The one million arXiv articles were divided into 20 periods in this part in order to increase the number of articles per period and reduce estimation error, which is not the same as the previous part. The identifier numbers of the first and last arXiv articles corresponding to each period are given in the section A of the Appendix.
The emergence of other LLMs is also inspired or influenced by ChatGPT, and we also assume that other LLMs have similar but not identical word preferences to ChatGPT. The most recent articles we processed were submitted in January 2024, when other LLMs should be less utilized. Therefore, we used ChatGPT for simulations.
@@ 3.2 LLM simulations
where $q^{d}_{ij}$ represents the word frequency of real abstracts in the dataset and $\tilde{q}^{d}_{ij}$ means the frequency after ChatGPT processing.
What are the prompts used in the real cases are still unclear, and we think simple prompts could better reflect the inherent word preferences of ChatGPT, as complex prompts may bring more human interference. So some simple prompts were used to reduce the bias due to prompts, for example,
@@
GPT-3.5 was utilized in our simulations for 10,000 abstracts in period 14 (April 2022 to July 2022), although it may not have the same word preferences as other LLMs. Many words have different frequencies before and after ChatGPT processing, such as the words “is”, “are”, and “significant” that we mentioned earlier. For simplicity, the results of the 4 categories with the highest number of articles are shown in Table 1 and the rest parts in this paper, namely cs (computer science), math (mathematics), astro (astrophysics), and cond-mat (condensed matter).
Table 1: Word frequency (per abstract) before and after ChatGPT processing.
words category before after change rate
is, are cs 2.01, 1.00 1.73, 0.83 -14%, -17%
is, are math 1.78, 0.74 1.61, 0.71 -9%, -5%
is, are astro 2.13, 1.39 1.90, 1.25 -11%, -1%
is, are cond-mat 2.00, 0.92 1.68, 0.80 -16%, -13%
@@
This corroborates the hypothesis, formulated earlier, that the drop in the frequency of these two words observed in real abstracts in 2023 may have been caused by ChatGPT. Combined with Figure 6 in the Appendix showing the correlation between changes in simulated and real data, we speculate that ChatGPT is one of the important reasons, possibly even the main reason, for the recent word frequency change in abstracts.
Our next step is to start by modeling LLM impact or ChatGPT impact, as well as estimating the impact based on real data and simulations. In order to minimize the impact of the research topic, different words should be used for estimation for different paper categories. And it is important to consider not only words that increase in frequency, but also those that decrease in frequency.
4 LLM impact
4.1 Simple model
Imagine different scenarios of using LLMs in scientific writing: a researcher might simply use it to correct grammatical errors, another employs it for translating native sentences into English, and yet another one wants it to polish their draft in English very purposefully. In theory, each of these use cases contributes the same proportion of LLM usage. But, as is well known, different prompts will lead to different outputs, which means different word frequency changes. Therefore, we use the more neutral term “LLM impact” instead of “proportion” in our estimation part. Because the estimates in this paper are based on ChatGPT simulations, it can also be called “ChatGPT impact”.
@@ 4.1 Simple model
$\bar{f}_{ij}(t)=(1-\eta_{j}(t))f^{*}_{ij}(t)+\eta_{j}(t)f^{*}_{ij}(t)(\bar{r}_{ij}+1)=f^{*}_{ij}(t)+\eta_{j}(t)f^{*}_{ij}(t)\bar{r}_{ij}$ (3)
where $\eta_{j}(t)$ denotes the proportion of abstracts in category $j$ affected by LLMs, and $f^{*}_{ij}(t)$ represents the original evolution in word frequency without LLMs.
Unfortunately, we cannot know the true value of $f^{*}_{ij}(t)$ in the LLM era, but we can replace it with the estimation $\hat{f}^{*}_{ij}(t)$ based on the word frequency before LLM was introduced. As our objective is to identify the words that LLM “likes” (or “dislikes”) to use compared to academic researchers on average, we assume that the frequencies of these words should remain stable without LLM, i.e., we take the average of the pre-ChatGPT periods before $T_{0}$ as following:
For a specified word $i$, we will have one estimate of $\eta_{j}(t)$, as $\bar{r}_{ij}$ and $f^{*}_{ij}(t)$ could be approximated with Eq. (1) and Eq. (4). We are also likely to get better results after combining the estimates of different words.
However, this model is highly idealized: we have to additionally consider the effects of noise (such as randomness inside LLM), uncertainty in word usage evolution without LLM, and the epistemic uncertainty in how users actually prompt LLMs.
@@ 4.2 Noise model
We now consider the noise terms, which might be modeled in many different ways.
For instance, we denote the word frequency for word $i$ in category $j$ by ${f}^{d}_{ij}$, which represents the word frequency observed in the data:
@@
where $\delta_{ij}(\cdot)$ represents noise and word usage variability which are not directly related to the internal parameters of LLM.
After taking into account the impact of LLMs, we split the word frequencies $f^{d}_{ij}(t)$ into two parts, $f^{\delta,\eta}_{ij}(t)$ and $f^{\delta,1-\eta}_{ij}(t)$, while they both have corresponding noise terms:
$f^{\delta,\eta}_{ij}(t)$ $=\eta_{j}(t)f^{*}_{ij}(t)+\delta_{ij}(\eta_{j}(t)f^{*}_{ij}(t))$ (6)
$f^{\delta,1-\eta}_{ij}(t)$ $=(1-\eta_{j}(t))f^{*}_{ij}(t)+\delta_{ij}((1-\eta_{j}(t))f^{*}_{ij}(t))\,.$ (7)
@@
$f^{d}_{ij}(t)=(1-\eta_{j}(t))f^{*}_{ij}(t)+\delta_{ij}((1-\eta_{j}(t))f^{*}_{ij}(t))+\mathrm{C}_{ij}(f^{\delta,\eta}_{ij}(t))=f^{\delta,1-\eta}_{ij}(t)+\mathrm{C}_{ij}(f^{\delta,\eta}_{ij}(t))$ (8)
where the function $\mathrm{C}_{ij}(\cdot)$ means the frequency after LLM process.
We assume that the noise in the “real” data and in the simulations due to LLM processing can be represented as $\epsilon_{ij}(\cdot)$ and $\epsilon_{ij}^{s}(\cdot)$, then Eq. (1) and Eq. (2) are related by
@@
Then, Eq. (8) – representing the difference in word frequency before and after LLM processing – can be rewritten as
@@
$\xi_{ij}(t)=$ $(\hat{r}_{ij}+1+\epsilon^{\eta}_{ij}(q,f,t))\delta_{ij}(\eta_{j}(t)f^{*}_{ij}(t))+\delta^{\prime}_{ij}((1-\eta_{j}(t))f^{*}_{ij}(t))\,.$ (15)
where $\delta^{\prime}_{ij}(\cdot)$ follows the same distribution as $\delta_{ij}(\cdot)$. It should be noted that $g_{ij}(t)$ includes only LLM-related noise $\epsilon_{ij}(\cdot)$ and $\epsilon^{s}_{ij}(\cdot)$, however $\xi_{ij}(t)$ contains $\delta_{ij}(\cdot)$ and $\delta^{\prime}_{ij}(\cdot)$ that are unrelated to LLM.